The following information applies to Division X and Division Y of Kent Corporation:



  



Chapter 24 Homework Question 7: (0.5 point) Required: Calculate the residual income for Division X. When entering the answers in Blackboard, omit $ signs. Use the minus sign for negative amount.

Chapter 24 Homework Question 8: (0.5 point) Required: Calculate the residual income for Division Y. When entering the answers in Blackboard, omit $ signs. Use the minus sign for negative amount.

Invested Assets
Inc. from Operations
Min. Rate Of Return Division X $800,000 $150,000 12.0% Division Y $600,000 $80,000 14.0%
Answer #1

Since it is not quite clear from the question how the data is plotted, just for clarity purposes, I have plotted the data below on the basis of which the question is answered :

Particulars Division X Division Y
Invested Assets 800,000$ 600,000$
Income from Operations 150,000$ 80,000$
Minimum Rate of Return 12% 14%

Answer to Question 7 :

Formula for calculating residual income is as follows :

Residual Income = Operating Income - (Invested Assets * Minimum rate of return required)

Hence, Residual income for Division X = 150,000 - (800,000*12%) = 150,000-96,000 = 54,000

Answer to Question 8 :

On the basis of formula mentioned above, residual income for division Y has been calculated below :

Residual Income for Division Y = 80,000 - (600,000*14%) = 80,000-84,000 = -4,000